OpenStudy (anonymous):

For an object falling freely from rest (disregarding air resistance), the distance the object falls varies directly as the square of the time. If an object is dropped from a 637 foot cliff and hits the ground in 7 seconds, how far did it fall in the first 3 seconds? 117 feet 273 feet 108 feet 137 feet

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (matlee):

137

OpenStudy (studygurl14):

"The distance the object falls varies directly as the square of the time" Distance: d time: t k = constant \(\large d=kt^2\) First, solve for k by plugging in 7 for t, and 637 for d. Next, find the answer by plugging the value you get for k back into the equation, and substitute 637 for d, and solve for t.

OpenStudy (studygurl14):

oops, sorry I mean for the next step, plug in 3 for t, and solve for d

OpenStudy (studygurl14):

Just do this first...lol: \(\large (637)=k(7)^2\) solve for k

OpenStudy (anonymous):

4459

OpenStudy (studygurl14):

nad by the way, @matlee 's answer is incorrect

OpenStudy (matlee):

dont listen to her i have a 50% in algerbra 2

OpenStudy (studygurl14):

nope @judtpleasegahd \(\large 637 = 49k\rightarrow k = \frac{637}{49}\tt =the~answer\)

OpenStudy (studygurl14):

by "the answer" I don't mean the answer of the entire problem. I just mean the value of k.

OpenStudy (anonymous):

13

OpenStudy (studygurl14):

and @matlee ....? That's not a good grade...

OpenStudy (studygurl14):

correct

OpenStudy (studygurl14):

Now, substitute 3 for t, and 13 for k in this equation, and solve for d \(\large d=kt^2\)

OpenStudy (matlee):

i think its pretty decent

OpenStudy (anonymous):

ty, not to be rude or mean at all but what do you mean by 50% ? @matlee

OpenStudy (studygurl14):

50% is an F

OpenStudy (matlee):

Like failing

OpenStudy (anonymous):

thats what i was thinking i just didnt want to be rude

OpenStudy (studygurl14):

In my mind, F is horrible, D is bad, C is ok, B is good, A is excellent

OpenStudy (matlee):

ty jud youve learened the lesson i wanted people to learn, ask before assuming :0 even if your asnwers could never be wrong just like math :p

OpenStudy (anonymous):

@StudyGurl14 same

OpenStudy (studygurl14):

so...how is it "pretty decent" @matlee ???

OpenStudy (anonymous):

let me stay focus lol @StudyGurl14 what were we saying

OpenStudy (studygurl14):

lol ok @judtpleasegahd

OpenStudy (matlee):

Lol goodluck to you guys goodluck onr your homewokrs

OpenStudy (studygurl14):

Let me type it again so you can see it clearly...@jud

OpenStudy (anonymous):

ok

OpenStudy (studygurl14):

Now, substitute 3 for t, and 13 for k in this equation, and solve for d \(\large d=kt^2\)

OpenStudy (anonymous):

3*13

OpenStudy (anonymous):

39?

OpenStudy (studygurl14):

not exactly...remember the exponent So... \(\large d=(13)(3)^2\)

OpenStudy (studygurl14):

And remember, \((3)^2=(3)(3)\)

OpenStudy (anonymous):

177

OpenStudy (studygurl14):

If you meant 117, then yes

OpenStudy (anonymous):

thank you :))

OpenStudy (studygurl14):

Anytime, my friend. \(\large\ddot\smile\)